# ChipletPart: Cost-Aware Partitioning for 2.5D Systems

### v2

@@ 6.3. Runtime remarks
ChipletPart is implemented entirely in C++, including a high-quality C++ translation of the Python-based cost model from (costmodelUCLA, ). Our optimized C++ translation achieves approximately a ${5}\times$ speedup compared to the original Python version. Empirical results indicate that ChipletPart, when integrated with the translated C++ cost model, attains over a $100\times$ speedup relative to an integration based on the Python cost model. Additionally, our open-sourced implementation is fully parallelized using OpenMP. These improvements make ChipletPart readily adaptable for any chiplet-based design flow. We present a detailed runtime breakdown of ChipletPart in Table 7. Approximately 67% of the execution time is spent during floorplanning, while 28% is spent on chiplet cost evaluation. Further details on ChipletPart’s runtime are provided in Table 6.
Scalability. For large designs such as WS4, our runtime remains under two hours. To improve scalability further, hierarchical decomposition techniques that exploit structural regularity in the netlist, along with aggressive solution pruning and potential usage of surrogate models for faster evaluation, can be employed.
@@ 6.4. Exploring Bayesian Optimization
Bayesian optimization (BO) is a powerful framework for optimizing expensive, black-box functions, particularly when gradients are unavailable and each function evaluation incurs significant computational overhead (shahriari2015taking, ). Unlike exhaustive or random search strategies, BO constructs a probabilistic surrogate model of the objective function and uses an acquisition function (gan2021acquisition, ) to balance exploration and exploitation. These properties make BO especially suitable for problems with complex constraints, cost landscape, and costly evaluations. To provide another quality-runtime tradeoff point, we have also explored the use of BO for chiplet partitioning. We jointly optimize (i) the assignment of IP blocks to chiplets, and (ii) the selection of technology nodes per chiplet to minimize total system cost. Directly optimizing the partitioning function $\phi:\mathcal{S}\rightarrow\mathcal{C}$ is infeasible due to the combinatorial explosion in the search space: for $k$ chiplets and $|\mathcal{S}|$ blocks, there are $k^{|\mathcal{S}|}$ possible assignments. Such a high-dimensional, discrete space is ill-suited to BO, which relies on continuous, low-dimensional representations (papenmeier2025, ). To address this, we use spectral embeddings to derive a continuous relaxation of the partitioning problem (chung1997spectral, ). We construct a normalized Laplacian matrix (merris1994laplacian, ) from the netlist connectivity graph and compute its $d$ nontrivial eigenvectors. These eigenvectors define a mapping:
@@
Formally, $x=[k,\mathbf{c}_{1},\ldots,\mathbf{c}_{k},\mathbf{t}_{1},\ldots,\mathbf{t}_{k}]$ and the dimensionality of $x$ is: $\dim(x)=1+k\cdot d+k\cdot m$. The BO engine constructs a surrogate model $\hat{\Phi}(x)$ to approximate the true cost $\Phi(x)$ and uses an acquisition function $\alpha(x)$ to guide exploration. We use a Gaussian Process (GP) (frazier2018tutorial, ) surrogate for small datasets (<50<50 IP blocks) and Random Forest (zhang2021prediction, ) for larger datasets. Our acquisition function comprises both Expected Improvement (gan2021acquisition, ) and Upper Confidence Bound (gan2021acquisition, ) metrics. Once BO picks a candidate $x$, it undergoes the following evaluation pipeline:
@@ 7. Additional Case Studies
We investigate how different technology parameters affect ChipletPart’s solutions. To accommodate an arbitrarily complex cost model, ChipletPart employs a black-box partitioning method rather than a mathematical programming approach. This design choice enables a broader exploration of the parameter space compared to other partitioners. To illustrate these benefits, we present case studies that examine effects of altering cost model parameters.
@@
Figure 13. (a) Reach limit vs. #chiplets. The partitioning solutions are generated with ChipletPart using reach values of $2\text{mm}$, $10\text{mm}$ and $20\text{mm}$. (b) IO type vs. #chiplets.
[…]
@@
IO type and reach. Chiplet systems often use reduced-size IO cells that trade driver strength and ESD protection for a smaller form factor. However, limited IO reach makes it challenging to generate feasible solutions. A $2\text{mm}$ reach may only allow direct connections between adjacent chiplets, while a larger reach ($20\text{mm}$) increases flexibility but also adds to system cost by increasing chiplet area. To explore this, we: (i) analyze the impact of IO reach on chiplet partitioning for a parallel IO testcase; and (ii) compare it with UCIe advanced and UCIe standard definitions (UCIe1p0, ), which have different cell sizes and reach. The parallel IO has $2\text{mm}$ reach, while the UCIe cases have $10\text{mm}$ reach.
Figure 13(a) shows that increasing reach enables a greater number of chiplets in the partitioning solution, while a short $2\text{mm}$ reach forces more blocks to remain together in fewer chiplets. This can be seen by looking at $TC_{2}$. Figure 13(b) compares the same parallel IO testcase with UCIe interfaces, showing that UCIe’s large IO cells favor fewer chiplets, while the parallel interface with smaller IOs and bump pitches enables finer partitioning.
@@
Mixed cost/power objective function. ChipletPart can simultaneously optimize cost and power using a weighted sum of chiplet cost (Equation 1) and power as its objective function. Table 9 shows results for various cost and power coefficient combinations. For this experiment, we use the WS1 and TC1 testcases, with all chiplets implemented in $7\text{nm}$ technology. We see that increasing the power coefficient leads to fewer chiplets as large IO drivers for inter-chiplet connections have higher power.
@@ 8. Conclusion
We have proposed ChipletPart, a novel chiplet partitioning framework designed to tackle the challenges of partitioning large, 2.5D heterogeneously integrated systems into chiplets. ChipletPart leverages an advanced cost model, is power-aware, and uses a genetic algorithm to simultaneously assign technologies to chiplets, all while consistently returning IO-feasible solutions that honor the driving reach of inter-chiplet IO drivers. Experimental results demonstrate that ChipletPart outperforms classical netlist partitioners, FloorPlet (ChenLZZL23, ), and human baselines. Additional studies illuminate dependencies of partitioning outcomes on the underlying packaging and IO technology. We have also explored Bayesian optimization as an alternative search strategy. While BO can sometimes yield better chiplet solutions compared to our genetic algorithm, it incurs significantly higher runtime overhead. Both GA and BO provide a choice of quality-runtime trade-offs.
While ChipletPart is designed as a cost-driven partitioning framework, its modular, black-box optimization approach promotes future integration of more complex objective functions that consider power, performance and thermal factors. Extending the current objective function to include additional factors remains a direction for future work. Considering detailed IO planning is also another direction for future work. Detailed IO planning that supports pass-through connections, multiple IO types on a die, routing congestion, use of “gas station” buffers (CoskunEJKMNS20, ), buffer chiplets, or other similar architectural considerations will bring significantly increased problem complexity. We also aim to enhance ChipletPart by identifying repeated structures in the netlist, a capability that could significantly improve scalability for larger problems. Finally, we are exploring the use of alternate optimizers, such as ILP, RL or BO. We also plan to integrate ChipletPart with OpenROAD (OpenROAD, ) to promote engagement in benchmarking and further development.

### v3

@@ 6.3. Runtime remarks
ChipletPart is implemented entirely in C++, including a high-quality C++ translation of the Python-based cost model from (costmodelUCLA, ). Our optimized C++ translation achieves approximately a ${5}\times$ speedup compared to the original Python version. Empirical results indicate that ChipletPart, when integrated with the translated C++ cost model, attains over a $100\times$ speedup relative to an integration based on the Python cost model. Additionally, our open-sourced implementation is fully parallelized using OpenMP (openmp, ). These improvements make ChipletPart readily adaptable for any chiplet-based design flow. We present a detailed runtime breakdown of ChipletPart in Table 8. Approximately 67% of the execution time is spent on floorplanning, while 28% is spent on chiplet cost evaluation. Further details on ChipletPart’s runtime are provided in Table 7.
Scalability. For large designs such as $WS_{4}$, our runtime remains under two hours. To improve scalability further, hierarchical decomposition techniques that exploit structural regularity in the netlist, along with aggressive solution pruning and potential usage of surrogate models for faster evaluation, can be employed.
@@ 6.4. Exploring Bayesian Optimization
Bayesian optimization (BO) is a powerful framework for optimizing expensive, black-box functions, particularly when gradients are unavailable and each function evaluation incurs significant computational overhead (shahriari2015taking, ). Unlike exhaustive or random search strategies, BO constructs a probabilistic surrogate model of the objective function and uses an acquisition function (gan2021acquisition, ) to balance exploration and exploitation. These properties make BO especially suitable for problems with complex constraints, cost landscape, and costly evaluations. To provide another quality-runtime tradeoff point, we have also explored the potential use of BO for chiplet partitioning. We jointly optimize (i) the assignment of IP blocks to chiplets, and (ii) the selection of technology nodes per chiplet to minimize total system cost. Directly optimizing the partitioning function $\phi:\mathcal{S}\rightarrow\mathcal{C}$ is infeasible due to the combinatorial explosion in the search space: for $k$ chiplets and $|\mathcal{S}|$ blocks, there are $k^{|\mathcal{S}|}$ possible assignments. Such a high-dimensional, discrete space is ill-suited to BO, which relies on continuous, low-dimensional representations (papenmeier2025, ). To address this, we use spectral embeddings to derive a continuous relaxation of the partitioning problem (chung1997spectral, ). We construct a normalized Laplacian matrix (merris1994laplacian, ) from the netlist connectivity graph and compute its $d$ nontrivial eigenvectors. These eigenvectors define a mapping:
@@
Formally, $x=[k,\mathbf{c}_{1},\ldots,\mathbf{c}_{k},\mathbf{t}_{1},\ldots,\mathbf{t}_{k}]$ and the dimensionality of $x$ is: $\dim(x)=1+k\cdot d+k\cdot m$. The BO engine constructs a surrogate model $\hat{\Phi}(x)$ to approximate the true cost $\Phi(x)$ and uses an acquisition function $\alpha(x)$ to guide exploration. We use a Gaussian Process (GP) (frazier2018tutorial, ) surrogate for small datasets (<50<50 IP blocks) and Random Forest (zhang2021prediction, ) for larger datasets. Our acquisition function comprises Expected Improvement (gan2021acquisition, ) and Upper Confidence Bound (gan2021acquisition, ) metrics. Once BO picks a candidate $x$, it undergoes the following evaluation pipeline:
@@
BO vs. GA trade-offs. Our results show that ChipletPart-GA provides the best overall quality/runtime tradeoff for chiplet partitioning: it converges quickly, is robust across benchmarks, and integrates naturally into a broader CAD flow. In contrast, ChipletPart-BO incurs significantly higher runtime due to surrogate-model construction and acquisition-function optimization, but can yield modestly better solutions on benchmarks whose spectral embeddings exhibit a smoother cost landscape (e.g., $WS_{3}$, $WS_{4}$). We recommend ChipletPart-GA for routine use or time-constrained flows, while ChipletPart-BO can potentially serve as a compute-intensive alternative that may improve quality on select, more challenging testcases.
@@ 7. Additional Case Studies
We investigate how different technology parameters affect ChipletPart’s solutions. To accommodate an arbitrarily complex cost model, ChipletPart employs a black-box partitioning method rather than a mathematical programming approach. This design choice enables a broader exploration of the parameter space compared to other partitioners. To illustrate these benefits, we present case studies that examine effects of altering the chiplet cost model parameters.
Image containing two bar charts showing different testcases with varied reach constraints and I/O types.
@@
Figure 14. (a) Reach limit vs. #chiplets. The partitioning solutions are generated with ChipletPart using reach values of 2 mm, 10 mm, and 20 mm. (b) I/O type vs. #chiplets.
@@
I/O type and reach. Chiplet systems often use reduced-size I/O cells that trade driver strength and electrostatic discharge (ESD) protection for a smaller form factor. However, limited I/O reach makes it challenging to generate feasible solutions. A 2 mm reach may only allow direct connections between adjacent chiplets, while a larger reach (e.g., 20 mm) increases flexibility but also adds to system cost by increasing chiplet area. To explore this, we: (i) analyze the impact of I/O reach on chiplet partitioning for a parallel I/O testcase; and (ii) compare the parallel I/O testcase with UCIe advanced and UCIe standard definitions (UCIe1p0, ), which have different cell sizes and reach. The parallel I/O has 2 mm reach, while the UCIe cases have 10 mm reach.
Figure 14(a) shows that increasing reach enables a greater number of chiplets in the partitioning solution, while a short 2 mm reach forces more blocks to remain together in fewer chiplets. As shown in Figure 4, longer I/O reach values enable connections that are impossible under shorter reach limits. A reach of 2 mm effectively restricts connectivity to nearest neighbors, which forces more blocks to remain colocated and increases overall cost. This effect is clearly illustrated in the $TC_{2}$ testcase: with a 2 mm reach, the resulting solution has a cost of 46.3, whereas increasing the reach to 10 mm produces a lower-cost solution (cost 43.4), corresponding to a 6.3% reduction. Figure 14(b) compares the same parallel I/O testcase with UCIe interfaces, showing that UCIe’s large I/O cells favor fewer chiplets, while the parallel interface with smaller I/Os and bump pitches enables finer partitioning. For example, this effect is evident in the $EPYC$ testcase. We verified the impact of I/O type and bump pitch by evaluating the cost of a fixed chiplet partitioning solution, the parallel_10um solution, under the other I/O configurations. When evaluated with the parallel_10um I/O configuration, the solution has a cost of 73.1, whereas evaluating the same partitioning solution under the UCIe_advanced_55um and UCIe_standard_110um configurations increases the cost to 74.1 and 74.5, respectively. Because this additional cost arises solely from the larger I/O cell sizes and pitches, the solutions optimized for UCIe_advanced_55um and UCIe_standard_110um reduce the total inter-chiplet connection bandwidth by merging blocks into fewer chiplets, thereby avoiding the higher I/O cost.
@@
Mixed cost/power objective function. ChipletPart can simultaneously optimize cost and power using a weighted sum of chiplet cost (Equation 1) and power as its objective function. Table 10 shows results for various cost and power coefficient combinations. For this experiment, we use the $WS_{1}$ and $TC_{1}$ testcases, with all chiplets implemented in 7 nm technology. As the power coefficient increases, the solutions favor fewer chiplets since large I/O drivers for inter-chiplet connections incur higher power consumption.
@@ 8. Conclusion
We have proposed ChipletPart, a novel chiplet partitioning framework designed to tackle the challenges of partitioning large, 2.5D heterogeneously integrated systems into chiplets. ChipletPart leverages an advanced cost model, is power-aware, and uses a genetic algorithm to simultaneously assign technologies to chiplets, all while consistently returning I/O-feasible solutions that honor the driving reach of inter-chiplet I/O drivers. Experimental results demonstrate that ChipletPart outperforms classical netlist partitioners, Floorplet (ChenLZZL23, ), and human baselines. Additional studies illuminate dependencies of partitioning outcomes on the underlying packaging and I/O technology. We have also explored Bayesian optimization as an alternative search strategy. Although Bayesian optimization can sometimes generate higher-quality solutions than the genetic algorithm, it incurs significantly higher runtime overhead.
While ChipletPart is designed as a cost-driven partitioning framework, its modular, black-box optimization approach promotes future integration of more complex objective functions that consider power, performance and thermal factors. Extending the current objective function to include additional factors remains a direction for future work. Considering detailed I/O planning is another direction for future work. Detailed I/O planning that supports pass-through connections, multiple I/O types on a die, “gas station” buffers (CoskunEJKMNS20, ), buffer chiplets, or other similar architectural considerations will bring significantly increased problem complexity. Evaluation with routing-congested designs is another topic of interest. We also aim to enhance ChipletPart by identifying repeated structures in the netlist, a capability that could significantly improve scalability for larger designs. Beyond 2.5D, adapting ChipletPart to 3D-stacked technologies, where vertical interconnects, TSV constraints, and tier locality introduce new partitioning objectives, is an opportunity for future work. Incorporating detailed thermal and power delivery models can potentially enable co-optimization for reliability, temperature, and system-level performance. Finally, we are exploring the use of alternate optimizers, such as Integer Linear Programming (ILP) and Reinforcement Learning (RL). We also plan to integrate ChipletPart with OpenROAD (OpenROAD, ) to promote engagement in benchmarking and further development.
Acknowledgements.
This work was supported in part by Samsung AI Center, Intel Corporation, DARPA/SRC CHIMES JUMP 2.0 center, NSF, and the CDEN center.